What is the percentage of p orbital character in diamond and graphite
 

Grade:12th pass

1 Answers

Arun
5 years ago

Graphite -

C atom is sp2 hybridised and thermodynamically more stable than diamond at ordinary temperature 

- wherein

High electrical conductivity is due to presence of delocalised  \pi -

 

Diamond -

C atom is sp3 hybridised and has three dimensional network structure

- wherein

It is bad conductor due to absence of valence electron

Sp^{2}

 
 
 
 
 
 hybridization 

 

%p = \frac{2}{3}\times 100

 
 
 
 
 

 

=  67%

Diamond 

Sp^{3}

 
 
 
 
 
 hybridisation 
%  p = \frac{3}{4}\times 100
 
 
 
 
 

 

= 75 %